DIRC Operations Report
Wednesday 2007/8/15 to Wednesday 2007/8/22
Oncall: Rolf Andreassen (Justine Serrano on backup)
Fast Monitoring: Rolf Andreassen, Justine Serrano
Luminosities this week
PEP peak lumi: ~1.1 x 1034
(last week: ~1.08 x 1034)
PEP delivery: 6.00 fb-1
(last week: 3.82 fb-1)
BaBar integrated lumi:5.68 fb-1
(last week: 3.68 fb-1)
Run 6 total delivery: 82.27 fb-1
Run 6 total BaBar recorded: 79.02 fb-1 total, including 7.89 fb-1 off-peak data
PEP activities/issues this week
08/16: Go back to on-resonance running.
BaBar activities/issues this week
DIRC activities/issues this week
08/15: Matt disabled the dump valve for the UPS battery replacement.
08/17: Inhibited the chiller SIAM, which had been glitching again.
08/22: As an experiment, tried to un-inhibit the chiller SIAM, since it seems to have been behaving nicely. It will get inhibited again if the glitches come back.
See DIRC logbook for details
Lumi and Background this month
- Take a standalone run at the next 1-2 hour down time to investigate the Xmas trees.
- Replace four DFBs with goups of 8 dead or flaky channels:
- sector 0, DFB 0.
- 2 in sector 1, DFB 1 and DFB 13.
- sector 3, DFB 6.
- sector 8, DFB 0.
- Unplug christmas tree in sector 5 and sector 11.
- Replace air filters on Wienner crates.
- Investigate the crate of sector 8 which has again killed the first 8 channels of DFB 7.
| Equipment || # used || # spare (*) || Comments |
| DFB || 168 || 18
6 good (3F,4E,64,66,68,69),
3 potentially good (6E,57,9C)
9 bad (05,70,6C,78,40,88,95,87,5A) in France
| DCC || 12 || 7
3 good (12,31,5)
2 bad (25,9) in France
- Justine and Rolf rotating backup & on-call commissioning duties.
- Justine leaving at the end of August, Loic will arrive in October.